OpenStudy (rebecca1588):

Eric has a gift card to his favorite restaurant. Every Tuesday he has the same meal at the restaurant. What is the equation that represents Eric’s gift card balance? Record any fractional answers in decimal form. _________ __________ (y – 100) =I________I (x – I_________I ) Eric’s Gift Card Weeks 2 4 6 8 Balance $125 $100 $75 $50

OpenStudy (rebecca1588):

(y-100)=12.5 (x-4) ?

OpenStudy (superdavesuper):

so the eqn given in the Q is: (y-100)=__(x-__) right?

OpenStudy (rebecca1588):

Yeah that's the question but I need someone to check my answer is this correct? (y-100)=12.5 (x-4) :)

OpenStudy (superdavesuper):

put x=6 n check for y...

OpenStudy (rebecca1588):

(y-100=-25/2 (x - 4) ?

OpenStudy (superdavesuper):

got it this time @rebecca1588 :)